# Comparison of Offset and Ratio Weighted Regressions in Tweedie Models with Application to Mid-Term Cancellations

### v1

@@ Abstract
In property and casualty insurance, particularly in automobile insurance, risk exposure is traditionally associated with the coverage duration. However, factors such as early contract cancellations demand more precise modelling to ensure accurate premium pricing. This study introduces and compares two approaches for modelling total claims (or loss costs) in insurance portfolios with a high proportion of policies that have partial-year exposure: the offset and ratio methods. We demonstrate that both approaches can be viewed as weighted regressions under the Tweedie distribution framework. Through an analysis based on the financial balance property, we find that the ratio approach outperforms the offset method. This comparison is illustrated using an automobile insurance portfolio, where a significant share of policyholders terminate their contracts before the coverage period concludes.
Keywords Tweedie distribution, weighted regression, offset variable, policy cancellation, stochastic order, positive definite matrix
@@ 1 Introduction
$\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right),$ (1.1)
where $\beta$ is a parameter vector to be estimated, and $\mathbf{X_{i}}=(x_{i,0},\dots,x_{i,q})^{T}$ represents the vector of risk characteristics associated with contract $i$ (with $q$ being a positive integer). These risk characteristics—such as age, sex, and the value of the insured goods—are used as covariates in the regression model, allowing the premium to vary based on these factors. The function $g(\cdot)$ serves as a link function that connects the score $\mathbf{X}_{i}^{T}\beta$ to the expected value, similar to the approach in Generalized Linear Models (GLMs) (see Nelder and Wedderburn (1972)).
@@
    The first approach assumes that the exposure $t_{i}$ is proportional to the mean of the conditional distribution $Y_{i}|\mathbf{X_{i}}$, where $f_{Y_{i}}(\cdot)$ represents the density or probability mass function of $Y_{i}|\mathbf{X_{i}}$, and $y_{i}$ denotes the observed loss cost for contract $i$. Under this assumption, the log-likelihood function for estimating the parameter vector $\beta$ is expressed as:
@@
    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right)\times t_{i}.$
@@
    $\exp\left(\mathbf{X}_{i}^{T}\beta\right)t_{i}=\exp\left(\mathbf{X}_{i}^{T}\beta+\log(t_{i})\right),$
@@
    To distinguish it from the offset approach, the mean of $Z_{i}|\mathbf{X_{i}}$ is assumed to no longer depend on $t_{i}$, and the density or probability mass function of $Z_{i}|\mathbf{X_{i}}$ incorporates the risk exposure through this weighting. The log-likelihood function for estimating $\beta$ is expressed as:
    $\ell(\bm{\beta}|z_{1},\ldots,z_{n})=\sum_{i=1}^{n}\log\left(f_{Z_{i}}(z_{i})\right),$
@@
    $\mathrm{E}\!\left[Z_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right),$
@@
    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=\mathrm{E}\!\left[t_{i}Z_{i}|\mathbf{X_{i}}\right]=\exp\left(\mathbf{X}_{i}^{T}\beta\right)\times t_{i}=\exp\left(\mathbf{X}_{i}^{T}\beta+\log(t_{i})\right),$
@@
It is important to emphasize that in the offset approach, the random variable $Z_{i}$ can be used to estimate the parameter vector $\beta$, but this method should not be confused with the ratio approach. Similarly, in the ratio approach, the random variable $Y_{i}$ can also be used, but this method should not ne mistaken for the offset approach.
@@ 1.1 Examples for claim counts
Under the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a Poisson distribution with mean $\zeta_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta})$. In this framework, it can be shown that the ratio approach corresponds to a Poisson weighted regression and yields the same inference results of the parameter vector $\beta$ as in the offset approach, where $Y_{i},i=1,\ldots,n$ is Poisson distributed with mean $\mu_{i}=t_{i}\exp(\mathbf{X}^{T}_{i}\bm{\beta})=\exp(\mathbf{X}^{T}_{i}\bm{\beta}+\log(t_{i}))$.
@@
    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+y_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+y_{i}\log(t_{i})-\log(y_{i}!)\right)$
    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{T}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)\right)\times t_{i}.$
@@
    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+t_{i}z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)-\log((t_{i}z_{i})!)\right)$
    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{T}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)\right)\times t_{i}.$
@@
In the offset approach, we assume that $Y_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\mu_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta}+\log(t_{i}))$ (from the Poisson distribution).
In the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\zeta_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta})$ (from the Poisson distribution).
It can be shown that these two approaches yield different inference results for the estimation of the parameter vector $\beta$.
@@
Depending on the distribution used for $Y_{i}$, it is not straightforward to determine whether both approaches are similar. Additionally, deciding which approach should be preferred in modelling is not simple. Even if the underlying distribution is the same, the parameters obtained from each approach may differ, which could result in varying premiums for each insured.
@@ 1.2 Structure of the paper
This paper emphasizes the differences between the offset and ratio regression approaches in modelling contract loss costs using the Tweedie distribution, rather than focusing on the number of claims. Section 2 introduces the notation and model assumptions, detailing the key properties of the Tweedie distribution and the parameter inference methods employed. Section 3 explains the offset and ratio approaches, as well as the link between them. We show that both approaches can be viewed as weighted regressions, where only the form of the weight differentiates the approaches. An analysis of asymptotic expectations, stochastic dominance, and the gap between aggregate loss costs and estimated premiums using simulated data is presented in Section 4. In Section 5, an empirical illustration of both approaches is provided, based on a sample of automobile insurance data from a Canadian insurer. The last section concludes.
@@ 2 GLM with Tweedie-distributed Response Variable
The loss cost of a contract is the total amount of claims, calculated as the aggregate sum of individual claim costs. Formally, let $N$ denote the he total number of claims for a contract, and let $Z_{k}$ represent the cost of the $k$-th claim, where $k,k=1,\ldots,N$ if $N>0$. Thus, assuming $Y$ denotes the loss cost of a contract, as $Y$ is obtained as follows:
@@
Even though individual claim costs $Z_{1},\ldots,Z_{N}$ are continuous random variables, we can demonstrate that the loss cost $Y$ is not a continuous random variable. Specifically, the probability of a null loss cost is non-zero, indicating that the loss cost has a mass at zero and is therefore a semi-continuous random variable. As a result, insurers often use the Tweedie distribution to model a contract premium when the loss cost is the response variable. The rationale for using the Tweedie distribution is its ability to represent a mixture distribution (both continuous and discrete), as noted by Tweedie et al. (1984). Additionally, Delong et al. (2021) provides a comprehensive summary of this mixture distribution, with relevant applications in actuarial science.
Accordingly, we consider the Tweedie distribution as a mixture distribution to model the premium using the loss cost as the response variable. The justification for employing the GLM theory in this context is based on the fact that the Tweedie distribution, as a mixture distribution, belongs to the linear exponential family of distributions, as discussed in Tweedie et al. (1984) and Jørgensen (1987). Before presenting the GLM approach in the context of the Tweedie distribution, we review the Tweedie distribution by referring to the works of Tweedie et al. (1984), Jørgensen (1987), and Delong et al. (2021).
@@ 2.1 Exponential Family Form of Tweedie Distribution
It is important to note that assuming the variance parameter lies within the interval ]1,2[]1,2[] 1 , 2 [ ensures that the Tweedie distribution density is defined as described above. As noted by Tweedie et al. (1984) and Jørgensen (1987), the Tweedie distribution can be generalized for any positive variance parameter $p>0$. However, they also mention that when the variance parameter lies within the interval ]0,1[]0,1[] 0 , 1 [, the Tweedie distribution does not belong to the exponential family. The density function $f(\cdot)$ mentioned above is expressed in the form of the exponential family density function. In the case of the Tweedie distribution, this family is sometimes referred to as the Exponential Dispersion Family (EDF) due to the inclusion of the dispersion parameter (see Jorgensen (1997)).
It should also be noted that $f(\cdot)$ lacks a specific form because the function $a(\cdot)$ is unknown. To determine $a(.)$, one can refer to Delong et al. (2021), which provides a comprehensive overview of the Tweedie distribution. In this paper, we do not focus on determining $a(\cdot)$ since it does not affect the regression approaches discussed here. Instead, we concentrate on the canonical parameter and the canonical link, which are relevant to insurance pricing. These two quantities for the Tweedie distribution are defined as follows:
@@
For the remainder of this paper, we denote the parameters of the Tweedie distribution by $(\mu,w,\phi,p)$ and we define $\mathcal{T}$ as the set of Tweedie distributions $(\mu,w,\phi,p)$ where the variance parameter $p$ lies within the interval ]1,2[]1,2[] 1 , 2 [.
@@ 2.2 Notations
  • $\mathbf{X}_{i}$: The vector of risk characteristics associated with contract $i$, defined as $\mathbf{X}_{i}=\left(1,x_{i,1},\ldots,x_{i,q}\right)^{T}$.
@@ 2.3 Assumptions and properties
    $Y_{i_{1}}|\mathbf{X}_{i_{1}}\perp\!\!\!\perp Y_{i_{2}}|\mathbf{X}_{i_{2}},\forall i_{1},i_{2}=1,\ldots,n\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ i_{1}\neq i_{2}$.
@@
The concept of GLM revolves around statistical inference on $\mu_{i}$, which is considered the premium for contract $i$ premium in insurance pricing. This consideration is justified by the fact that the mean parameter $\mu_{i}$, regardless of the distribution of $Y_{i}|\mathbf{X}_{i}$, corresponds to the constant closest to $Y_{i}|\mathbf{X}_{i}$ in terms of Euclidean distance (for further details, see Denuit et al. (2007)).
@@
$\mu_{i}$ $=\exp\left(\bm{X}^{T}_{i}\beta\right)=\exp\left(\beta_{0}\right)\prod_{j=1}^{q}\exp\left(\beta_{j}x_{i,j}\right)=b_{0}\times\prod_{j=1}^{q}r_{j},$ (2.2)
@@
  • $\beta$: The parameter vector, defined as $\beta=(\beta_{0},\beta_{1},\ldots,\beta_{q})^{T}$ where $\beta_{j}\in\mathbf{R}$ for $j=0,1,\ldots,q$;
@@
The advantage of using the logarithm function as the link function is that the premium $\mu_{i}$ for contract $i$ can be expressed as the product of the basic premium $b_{0}$ and relativity factors $r_{j},j=1,\ldots,q$. While the basic premium $b_{0}$ is consistent across all contracts, the relativity factors depend on the specific risk characteristics of each contract. Consequently, statistical inference on $\mu_{i}$ is equivalent to estimating the parameter vector $\beta$.
@@
If a random variable $Z$ is Tweedie-distributed with parameters $(\mu,w,\phi,p)$ and belongs to the set $\mathcal{T}$, then $tZ$, for any positive $t$, is also Tweedie-distributed with parameters $(t\mu,\frac{w}{t^{2-p}},\phi,p)$ and belongs to the same set $\mathcal{T}$. Conversely, if $tZ$, for any positive $t$, is Tweedie distributed with parameters $(\mu,w,\phi,p)$ and belongs to the set $\mathcal{T}$, then $Z$ is also Tweedie-distributed with parameters$(\frac{\mu}{t},wt^{2-p},\phi,p)$ and belongs to the same set $\mathcal{T}$. Formally, we have:

### v2

@@ Abstract
In property and casualty insurance, particularly in automobile insurance, risk exposure is commonly assumed to be proportional to the duration of coverage. This assumption leads to two standard estimation strategies: the ratio approach, which normalizes the response variable (e.g., claim cost or premium) by the exposure, and the offset approach, which incorporates a transformation of the exposure—typically its logarithm—as a fixed regressor in the mean structure of the model. Although both approaches rely on the same proportionality assumption, they are not equivalent when the response variable follows a Tweedie distribution, a framework widely used in insurance analytics. In this paper, we show that each approach can be implemented independently and yields a consistent estimator of the true mean parameter vector. We then show that the offset approach is asymptotically more efficient than the ratio approach, a result established both theoretically and through simulation studies. However, when evaluated from the perspective of portfolio-level financial balance, the ratio approach exhibits superior performance, particularly in the presence of heterogeneous or truncated exposures arising from mid-term policy cancellations. These theoretical results are illustrated through an empirical analysis of an automobile insurance portfolio with a high cancellation rate, highlighting the practical implications of model choice for premium estimation under variable exposure conditions.
Keywords Tweedie distribution, weighted regression, offset variable, policy cancellation, positive definite matrix
@@ 1 Introduction
$\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right),$ (1.1)
where $\bm{\beta}$ is a parameter vector to be estimated, and $\mathbf{X_{i}}=(x_{i,0},\dots,x_{i,q})^{\top}$ represents the vector of risk characteristics associated with contract $i$ (with $q$ being a positive integer). These risk characteristics—such as age, sex, and the value of the insured goods—are used as covariates in the regression model, allowing the premium to vary based on these factors. The function $g(\cdot)$ serves as a link function that connects the score $\mathbf{X}_{i}^{\top}\bm{\beta}$ to the expected value, similar to the approach in Generalized Linear Models (GLMs) (see Nelder and Wedderburn (1972)).
@@
    The first approach assumes that the exposure $t_{i}$ is proportional to the mean of the conditional distribution $Y_{i}|\mathbf{X_{i}}$, where $f_{Y_{i}}(\cdot)$ represents the density or probability mass function of $Y_{i}|\mathbf{X_{i}}$, and $y_{i}$ denotes the observed loss cost for contract $i$. Under this assumption, the log-likelihood function for estimating the parameter vector $\bm{\beta}$ is expressed as:
@@
    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right)\times t_{i}.$
@@
    $\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right)t_{i}=\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}+\log(t_{i})\right),$
@@
    To distinguish it from the offset approach, the mean of $Z_{i}|\mathbf{X_{i}}$ is assumed to no longer depend on $t_{i}$, and the density or probability mass function of $Z_{i}|\mathbf{X_{i}}$ incorporates the risk exposure through this weighting. The log-likelihood function for estimating $\bm{\beta}$ is expressed as:
    $\ell(\bm{\beta}|z_{1},\ldots,z_{n})=\sum_{i=1}^{n}t_{i}\log\left(f_{Z_{i}}(z_{i})\right),$
@@
    $\mathrm{E}\!\left[Z_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right),$
@@
    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=\mathrm{E}\!\left[t_{i}Z_{i}|\mathbf{X_{i}}\right]=\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right)\times t_{i}=\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}+\log(t_{i})\right),$
@@
It is important to emphasize that in the offset approach, the random variable $Z_{i}$ can be used to estimate the parameter vector $\bm{\beta}$, but this method should not be confused with the ratio approach. Similarly, in the ratio approach, the random variable $Y_{i}$ can also be used, but this method should not ne mistaken for the offset approach.
@@ 1.1 Examples for claim counts
Under the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a Poisson distribution with mean $\zeta_{i}=\exp(\mathbf{X}^{\top}_{i}\bm{\beta})$. In this framework, it can be shown that the ratio approach corresponds to a Poisson weighted regression and yields the same inference results of the parameter vector $\bm{\beta}$ as in the offset approach, where $Y_{i},i=1,\ldots,n$ is Poisson distributed with mean $\mu_{i}=t_{i}\exp(\mathbf{X}^{\top}_{i}\bm{\beta})=\exp(\mathbf{X}^{\top}_{i}\bm{\beta}+\log(t_{i}))$.
@@
    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)+y_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)+y_{i}\log(t_{i})-\log(y_{i}!)\right)$
    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{\top}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)\times t_{i}.$
@@
    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)+t_{i}z_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)-\log((t_{i}z_{i})!)\right)$
    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{\top}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)\times t_{i}.$
@@
In the offset approach, we assume that $Y_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\mu_{i}=\exp(\mathbf{X}^{\top}_{i}\bm{\beta}+\log(t_{i}))$ (from the Poisson distribution). In the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\zeta_{i}=\exp(\mathbf{X}^{\top}_{i}\bm{\beta})$ (from the Poisson distribution). It can be shown that these two approaches yield different inference results for the estimation of the parameter vector $\bm{\beta}$.
@@
Depending on the distribution used for $Y_{i}$, it is not straightforward to determine whether both approaches are equivalent. Even when the underlying distribution is assumed to be the same, the estimated parameters obtained from each approach may differ, potentially leading to different premiums for individual policyholders. Ideally, if the data-generating process for insurance claims were fully understood, one could determine which modelling approach is most appropriate. However, in practice, this is rarely the case, making it difficult to justify one formulation over the other on theoretical grounds alone. Consequently, the choice between approaches often relies on a combination of empirical performance and desirable mathematical properties.
In this context, it is worth noting that Denuit and Trufin (2019) already provide some comparisons between these two approaches. In particular, they show that the offset approach can be interpreted as a weighted regression when the response variable follows a Tweedie distribution. Building on their work, we extend the comparison further by emphasizing the role of stochastic orders and the principle of financial balance, which provide additional insights into the structural differences between the two modeling approaches.
@@ 1.2 Structure of the paper
This paper emphasizes the differences between the offset and ratio regression approaches in modeling contract loss costs using the Tweedie distribution, rather than focusing on claim frequency. Section 2 introduces the notation and model assumptions, detailing key properties of the Tweedie distribution and the parameter inference methods used. It also provides the technical framework required to establish the equivalence results and the asymptotic comparisons developed in the remainder of the paper. Section 3 then explains the offset and ratio approaches, as well as the link between them. We show that both approaches can be viewed as weighted regressions, where only the form of the weight differentiates the approaches. Section 4 presents an analysis of the consistency of the parameter vector estimators for both approaches, demonstrates the asymptotic efficiency of the offset-approach parameter vector estimator, and examines the gap between observed aggregate loss costs and the sum of estimated premiums based on simulated data. In Section 5, we provide an empirical illustration of both approaches using a sample of automobile insurance data from a Canadian insurer. Finally, Section 6 concludes the paper.
@@ 2 GLM with Tweedie-distributed Response Variable
The loss cost of a contract is the total amount of claims, calculated as the aggregate sum of individual claim costs. Formally, let $N$ denote the total number of claims for a contract, and let $Z_{k}$ represent the cost of the $k$-th claim, where $k,k=1,\ldots,N$ if $N>0$. Thus, assuming $Y$ denotes the loss cost of a contract, as $Y$ is obtained as follows:
@@
Even though individual claim costs $Z_{1},\ldots,Z_{N}$ are continuous random variables, we can demonstrate that the loss cost $Y$ is not a continuous random variable. Specifically, the probability of a null loss cost is non-zero, indicating that the loss cost has a mass at zero and is therefore a semi-continuous random variable. As a result, insurers often use the Tweedie distribution to model a contract premium when the loss cost is the response variable. The rationale for using the Tweedie distribution is its ability to represent a mixture distribution (both continuous and discrete), as noted by Tweedie and others (1984). Additionally, Delong et al. (2021) provides a comprehensive summary of this mixture distribution, with relevant applications in actuarial science.
Accordingly, we model the premium using the loss cost as the response variable, assuming that the latter follows a Tweedie distribution. This distribution is viewed as a compound Poisson–Gamma mixture and belongs to the linear exponential family, which justifies the use of the Generalized Linear Model (GLM) framework (see Tweedie, 1984; Jørgensen, 1987). Before introducing the GLM approach with the Tweedie distribution, we briefly review its properties by referring to Tweedie and others (1984), Jørgensen (1987), and Delong et al. (2021). It is also important to note that several results presented here can also be found in Wüthrich and Merz (2023) and Denuit and Trufin (2019); however, this review is essential for a proper understanding of the offset and ratio regression approaches.
@@ 2.1 Exponential Family Form of Tweedie Distribution
It is important to note that assuming the variance parameter lies within the interval ]1,2[]1,2[ ensures that the Tweedie distribution density is defined as described above. As noted by Tweedie and others (1984) and Jørgensen (1987), the Tweedie distribution can be generalized for any positive variance parameter $p>0$. However, they also mention that when the variance parameter lies within the interval ]0,1[]0,1[, the Tweedie distribution does not belong to the exponential family. The density function $f(\cdot)$ mentioned above is expressed in the form of the exponential family density function. In the case of the Tweedie distribution, this family is sometimes referred to as the Exponential Dispersion Family (EDF) due to the inclusion of the dispersion parameter (see Jorgensen (1997)).
It should also be noted that $f(\cdot)$ lacks a specific form because the function $a(\cdot)$ is unknown. To determine $a(\cdot)$, one can refer to Delong et al. (2021), which provides a comprehensive overview of the Tweedie distribution. In this paper, we do not focus on determining $a(\cdot)$ since it does not affect the regression approaches discussed here. Instead, we concentrate on the canonical parameter and the canonical link, which are relevant to insurance pricing. These two quantities for the Tweedie distribution are defined as follows:
@@
For the remainder of this paper, we denote the parameters of the Tweedie distribution by $(\mu,w,\phi,p)$ and we define $\mathcal{T}$ as the set of Tweedie distributions $(\mu,w,\phi,p)$ where the variance parameter $p$ lies within the interval ]1,2[]1,2[.
@@ 2.2 Notations
  • $\mathbf{X}_{i}$: The vector of risk characteristics associated with contract $i$, defined as $\mathbf{X}_{i}=\left(x_{i,0},x_{i,1},\ldots,x_{i,q}\right)^{\top}$, where $x_{i,0}=1,\forall i=1,\ldots,n$.
@@ 2.3 Assumptions and properties
    $Y_{i_{1}}|\mathbf{X}_{i_{1}}\perp\!\!\!\perp Y_{i_{2}}|\mathbf{X}_{i_{2}},\forall i_{1},i_{2}=1,\ldots,n\penalty 10000\ \text{such that}\penalty 10000\ i_{1}\neq i_{2}$.
@@
The concept of GLM revolves around statistical inference on $\mu_{i}$, which is considered the premium for contract $i$ in insurance pricing. This consideration is justified by the fact that the mean parameter $\mu_{i}$, regardless of the distribution of $Y_{i}|\mathbf{X}_{i}$, corresponds to the constant closest to $Y_{i}|\mathbf{X}_{i}$ in terms of Euclidean distance (for further details, see Denuit et al. (2007)).
@@
$\mu_{i}$ $=\exp\left(\bm{X}^{\top}_{i}\bm{\beta}\right)=\exp\left(\beta_{0}\right)\prod_{j=1}^{q}\exp\left(\beta_{j}x_{i,j}\right)=b_{0}\times\prod_{j=1}^{q}r_{j},$ (2.2)
@@
  • $\bm{\beta}$: The parameter vector, defined as $\bm{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{q})^{\top}$ where $\beta_{j}\in\mathbf{R}$ for $j=0,1,\ldots,q$;
@@
The advantage of using the logarithm function as the link function is that the premium $\mu_{i}$ for contract $i$ can be expressed as the product of the basic premium $b_{0}$ and relativity factors $r_{j},j=1,\ldots,q$. While the basic premium $b_{0}$ is consistent across all contracts, the relativity factors depend on the specific risk characteristics of each contract. Consequently, statistical inference on $\mu_{i}$ is equivalent to estimating the parameter vector $\bm{\beta}$.
@@
[…]